OpenStudy (anonymous):

#5 #6 #7 #8 #9 Help with math

OpenStudy (anonymous):

what kind ?

OpenStudy (anonymous):

Look at the picture

OpenStudy (anonymous):

#5 If 2t + 3 = -9 what is the value of -3t - 7? #6 Solve 2x - 4 = -7 justify each step #7 Your chorus holds a car wash they have $25.00 for making change at the end of the car wash they have $453.50 how much money did they make? #8 The rate to rent a certain truck is $55 per day and $0.20 per mile your family pays $80 to rent this truck for 1 day how many miles did your family drive? #9 Movie tickets for an adult and 3 children cost $20 and adult ticket cost $2 more than a child's ticket find the cost of the adult ticket?

OpenStudy (retireed):

#5 If 2t + 3 = -9 Solve for t 2t + 3 - 3 = -9 - 3 Subtract 3 from each side 2t = -12 Solve for t ---------------------------------------- Now Plug in the value of t into the second equation what is the value of -3t - 7? -3(t) - 7 = _________???

OpenStudy (anonymous):

t=10?

OpenStudy (retireed):

No 2t = -12 Divide both sides by 2

OpenStudy (anonymous):

oh sorry i did the wrong problem

OpenStudy (anonymous):

-6

OpenStudy (retireed):

2t = -12 Divide both sides by 2 2t /2 = -12 / 2 t = ___________???

OpenStudy (retireed):

Yes, -6

OpenStudy (anonymous):

can u help me with the rest?

OpenStudy (retireed):

Now Plug in the value of t into the second equation t = -6 what is the value of -3t - 7? -3(-6) - 7 = _________???

OpenStudy (anonymous):

11

OpenStudy (retireed):

Yes, 18 - 7 = 11

OpenStudy (retireed):

#6 Solve 2x - 4 = -7 justify each step Step 1) Isolate the 2x by adding 4 to both sides

OpenStudy (anonymous):

so the answer for #5 is 11

OpenStudy (retireed):

Yes, 18 - 7 = 11 #6 Solve 2x - 4 = -7 justify each step Step 1) Isolate the 2x by adding 4 to both sides

OpenStudy (anonymous):

so add 4 to both sides of the problem

OpenStudy (retireed):

#6 Solve 2x - 4 = -7 justify each step Step 1) Isolate the 2x by adding 4 to both sides Step 2) Isolate the x, Divide both sides by 2

OpenStudy (anonymous):

ok so 6x - 8 = -7?

OpenStudy (anonymous):

is the answer -3/2